# Holmes: Towards Distributed Training Across Clusters with Heterogeneous NIC Environment

### v3

@@ 5.1 Distributed Training Frameworks
Some popular distributed training frameworks tailored for LLMs include Megatron-LM [30, 23, 3], Megatron-LLaMA [2], and DeepSpeed [25, 28, 26, 1]. Megatron-LM, developed by NVIDIA, efficiently handles training of very large models using a combination of model parallelism and data parallelism across multiple GPUs and nodes. This approach enables LLM training by sharing model parameters among GPUs and processing data in parallel [23, 36]. Megatron-LM excels in minimizing synchronization delays during training through optimized communication between nodes and GPUs, utilizing efficient communication and synchronization techniques. Megatron-LLaMA, an extension of Megatron-LM, enhances the framework by integrating a standardized LLaMA implementation and introducing an efficient communication-computation parallelism method. Another notable framework is DeepSpeed, an open-source library from Microsoft designed for efficient training of large-scale deep learning models. DeepSpeed features memory-efficient training, accommodating models larger than GPU memory capacity through techniques like mixed-precision training, gradient accumulation, and offloading optimizer states. Its ZeRO optimization technology further improves memory efficiency by distributing model weights and optimizer states among GPUs and nodes, reducing memory usage and enabling training of even larger models [25, 28].
5.2 Communication Optimization
Distributed training necessitates frequent parameter synchronization. In the initial stages of communication scheduling, the primary focus is on ensuring model consistency [33, 27, 13] and refining parameter synchronization architectures [18, 24, 11]. With the increasing scale of models, certain research endeavors explore coordinated optimization of both communication and parallel strategies [34, 40]. This involves implementing advanced strategies to enhance parallel efficiency while concurrently reducing communication demands.
Recent advancements in distributed training efficiency, exemplified by TopoOpt [34], focus on coordinating improvements in computation, communication, and network topology. TopoOpt employs alternative optimization techniques and draws inspiration from group theory, utilizing the TotientPerms algorithm to optimize network topology, routing plans, and parallelization strategies. Another noteworthy project, Alpa [40], addresses challenges in cross-mesh sharding by introducing a broadcast-based communication mechanism and an overlap-friendly pipeline scheduling strategy.
6 Conclusion

### v4

@@ 5.1. Distributed Training Frameworks
Some popular distributed training frameworks tailored for LLMs include Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; NVIDIA, 2022a), Megatron-LLaMA (Alibaba, 2023), and DeepSpeed (Rajbhandari et al., 2020; Ren et al., 2021; Rasley et al., 2020; MicroSoft, 2023). Megatron-LM, developed by NVIDIA, efficiently handles training of very large models using a combination of model parallelism and data parallelism across multiple GPUs and nodes. This approach enables LLM training by sharing model parameters among GPUs and processing data in parallel (Narayanan et al., 2021b; Yi et al., 2017). Megatron-LM excels in minimizing synchronization delays during training through optimized communication between nodes and GPUs, utilizing efficient communication and synchronization techniques. Megatron-LLaMA, an extension of Megatron-LM, enhances the framework by integrating a standardized LLaMA implementation and introducing an efficient communication-computation parallelism method. Another notable framework is DeepSpeed, an open-source library from Microsoft designed for efficient training of large-scale deep learning models. DeepSpeed features memory-efficient training, accommodating models larger than GPU memory capacity through techniques like mixed-precision training, gradient accumulation, and offloading optimizer states. Its ZeRO optimization technology further improves memory efficiency by distributing model weights and optimizer states among GPUs and nodes, enabling training of even larger models (Rajbhandari et al., 2020; Ren et al., 2021).
5.2. Communication Optimization
Distributed training necessitates frequent parameter synchronization. In the initial stages of communication scheduling, the primary focus is on ensuring model consistency (Valiant, 1990; Recht et al., 2011; Ho et al., 2013) and refining parameter synchronization architectures (Li et al., 2014; Patarasuk and Yuan, 2009; Geng et al., 2018). With the increasing scale of models, certain research endeavors explore coordinated optimization of both communication and parallel strategies (Wang et al., 2023; Zhuang et al., 2023). This involves implementing advanced strategies to enhance parallel efficiency while concurrently reducing communication demands.
Recent advancements in distributed training efficiency, exemplified by TopoOpt (Wang et al., 2023), focus on coordinating improvements in computation, communication, and network topology. TopoOpt employs alternative optimization techniques and draws inspiration from group theory, utilizing the TotientPerms algorithm to optimize network topology, routing plans, and parallelization strategies. Another noteworthy project, Alpa (Zhuang et al., 2023), addresses challenges in cross-mesh sharding by introducing a broadcast-based communication mechanism and an overlap-friendly pipeline scheduling strategy.
6. Conclusion
@@ 7. Acknowledgement
APPENDIX A
ARTIFACT DESCRIPTION
We present the reproducibility artifact of our experimental validation, including the description of the computational artifacts. Our main contributions of the article lies in Holmes, an LLM training framework across GPU clusters with different type of RDMA NICs. The reproducibility artifact can assist in replicating the comparative effects mentioned in the paper with other models or methods, thereby demonstrating the efficacy of our work.
Artifact Dependencies and Requirements
  • Hardware resources: XEON-GPU, NVIDIA HGX, 200G Infiniband *4 or 200G ROCE *2.
  • Operating system: Ubuntu 18.04.4 LTS.
  • Software libraries Dependencies: Python-3.8.10, PyTorch-2.1.0, CUDA Version: 12.1.
  • Input dataset: we use OPT WebText dataset as input dataset to train GPT model.
  • Other dependencies or requirements: python packages including pybind11, transformers, nltk, etc.
Artifact Installation and Deployment Process
  • Our code base is available at https://anonymous.4open.science
    /r/Holmes-40A5
    . As for the runtime environment, we strongly recommend using the latest PyTorch, cuda, nccl and APEX release. You can launch an instance of the official PyTorch image or the runnable image provided by us and mount Holmes code, your dataset with the corresponding Docker commands. And install other dependencies according to the requirement file.
Reproducibility of Experiments
  • Our code base is capable of efficiently training large language models with both model and data parallelism over the heterogeneous NIC environment (including InfiniBand, RoCE and Ethernet). We’ve provided several scripts for pretraining GPT model in the examples directory of code base. To execute the script for evaluation, certain variables need to be set, including model scale(hidden size, number of heads and layers), heterogeneous NIC environment information (NUM_IB_BLOCK, NUM_GPUS_PER_IB_BLOCK), optimization training strategies switches(use-hetnet, use-asymmetric-pipeline-division, overlapped-distributed-optimizer), and NCCL network config parameters(SOCKET_IFNAME, HCA).
  • To demonstrate the framework scaling on multiple nodes and model sizes, we set variable hidden size, number of attention heads and layers to get specific model size and use up to 96 A100 GPUs (equipped with InfiniBand and RoCE NIC) to pretrain GPT models. The output results include metrics such as throughput, TFLOPS and communication operations overhead, which reflect the performance of our framework in training large language models.
  • To demonstrate that Holmes reduces communication overhead during training in heterogeneous environments, we compared the Grads-Reduce-Scatter-communication time with different parameter groups in both homogeneous and heterogeneous NIC environments. And through turning on the optimization switch in the script, we compare the performance of different parameter groups in training GPT model with Self-Adapting Pipeline Partition strategy and uniform partitioning method. Additionally, we conducted experiments on other mainstream training frameworks including Megatron-LM, Megatron-DeepSpeed, Megatron-LLaMa, and compared Holmes with these frameworks in terms of throughput, TFLOPS, and speedup rations when training the same parameterized large language models with the same number of nodes.